# Neural Spacetimes for DAG Representation Learning

### v1

@@ Abstract
We propose a class of trainable deep learning-based geometries called Neural Spacetimes (NSTs), which can universally represent nodes in weighted directed acyclic graphs (DAGs) as events in a spacetime manifold. While most works in the literature focus on undirected graph representation learning or causality embedding separately, our differentiable geometry can encode both graph edge weights in its spatial dimensions and causality in the form of edge directionality in its temporal dimensions. We use a product manifold that combines a quasi-metric (for space) and a partial order (for time). NSTs are implemented as three neural networks trained in an end-to-end manner: an embedding network, which learns to optimize the location of nodes as events in the spacetime manifold, and two other networks that optimize the space and time geometries in parallel, which we call a neural (quasi-)metric and a neural partial order, respectively. The latter two networks leverage recent ideas at the intersection of fractal geometry and deep learning to shape the geometry of the representation space in a data-driven fashion, unlike other works in the literature that use fixed spacetime manifolds such as Minkowski space or De Sitter space to embed DAGs. Our main theoretical guarantee is a universal embedding theorem, showing that any $k$-point DAG can be embedded into an NST with $1+\mathcal{O}(\log(k))$ distortion while exactly preserving its causal structure. The total number of parameters defining the NST is sub-cubic in $k$ and linear in the width of the DAG. If the DAG has a planar Hasse diagram, this is improved to $\mathcal{O}(\log(k))+2)$ spatial and 2 temporal dimensions. We validate our framework computationally with synthetic weighted DAGs and real-world network embeddings; in both cases, the NSTs achieve lower embedding distortions than their counterparts using fixed spacetime geometries.
@@ 1 Introduction
Graphs are a ubiquitous mathematical abstraction used in practically every branch of science from the analysis of social networks (Robins et al., 2007) to economic stability (Hurd et al., 2016) and genomics (Marbach et al., 2012). The breadth of social and physical phenomena encoded by large graphs has motivated the machine learning community to seek efficient representations (or embeddings) of graphs in continuous spaces. Since the structure of graphs typically has properties very different from those of standard Euclidean spaces (e.g., trees exhibit exponential volume growth), a significant effort has been made to identify non-Euclidean geometries which can faithfully capture the structure of general graphs. Examples include hyperbolic embeddings of trees and tree-like graphs (Ganea et al., 2018; Sonthalia & Gilbert, 2020; Shimizu et al., 2021; Kratsios et al., 2023a, b), spherical and toroidal embeddings of graphs with cycles (Schoenberg, 1942; Guella et al., 2016; Giovanni et al., 2022b), embeddings of graphs with several of these characteristics into product Riemannian geometries (Borde et al., 2023a, b; Giovanni et al., 2022a), alternative combinations of constant curvature Riemannian manifolds (Lu et al., 2023) and manifolds with locally controllable Ricci curvature (Giovanni et al., 2022b).
@@
Figure 1: A Neural Spacetime (NST) is a learnable triplet $\mathcal{S}=(\mathcal{E},\mathcal{D},\mathcal{T})$, where $\mathcal{E}:\mathbb{R}^{N}\rightarrow\mathbb{R}^{D+T}$ is a (feature) encoder network, $\mathcal{D}:\mathbb{R}^{D+T}\times\mathbb{R}^{D+T}\to[0,\infty)$ is a learnable quasi-metric on $\mathbb{R}^{D}$ and $\mathcal{T}:\mathbb{R}^{D+T}\to\mathbb{R}^{T}$ is a learnable partial order on $\mathbb{R}^{T}$. Given an input Directed Acyclic Graph (DAG), $\mathcal{E}$ optimizes the location of the nodes $u,v,w$ as events in the space-time manifold $\hat{u},\hat{v},\hat{w}$, while concurrently $\mathcal{D}$ and $\mathcal{T}$ learn the geometry of space and time themselves. The objective is to find a geometry that can faithfully represent, with minimal distortion, the metric geometry of the input DAG in space as well as its causal connectivity in time.
Despite these successes, it was shown in Borde & Kratsios (2023, Proposition 2) that no complete Riemannian geometry can contain an isometric copy of even very simple graph structures, highlighting the inherent limitation of such spaces in representation learning. In the same paper, the authors developed a trainable deep learning-based fractal geometry, which is flexible enough to exactly (isometrically) represent any undirected graph. In particular, the architecture avoided the need to manually identify the appropriate closed-form metric to embed different graphs and allowed for the optimization of a parametric metric via gradient descent instead (see Appendix A.3). Although distance information (e.g. shortest path graph geodesic distance) alone is enough to recover undirected edges, standard metric learning does not easily encode direction; thus far, the majority of works has focused on representation learning of undirected graphs. However, the description of many real-world systems requires directed graphs, for example, gene regulatory networks (Marbach et al., 2012), flow networks (Deleu et al., 2022), stochastic processes (Backhoff-Veraguas et al., 2020), or graph metanetworks (Lim et al., 2024) to name a few.
@@
The Directed Graph Embedding Problem. Directed graphs (or digraphs) play a significant role in causal reasoning, where they are used to study the relationship between a cause and its effect in a complex system. These problems make use of representation spaces with causal structures, modeling the directional edges of a discrete graph as causal connections in a continuous representation space (Zheng et al., 2018). However, most approaches in causal representation learning primarily focus on capturing directional information while neglecting distance information. Recent works (Law & Lucas, 2023; Sim et al., 2021) have suggested that Lorentzian spacetimes from general relativity are suitable representation spaces for directed graphs, as these geometries possess an arrow of time that provides them with causal structure (see Appendix A.5) while also being capable of encoding distance. In particular, some Lorentzian spacetimes called globally hyperbolic (Geroch, 1970) have the important property that any pair of causally related points may be joined by a causal geodesic segment of maximal length (Beem et al., 1996). It is important to note that these geometries do not allow for time loops, where points can have a non-trivial causal relationship with themselves in time. This means they cannot model general digraphs, only DAGs, which are also the focus of our work. Globally hyperbolic Lorentzian spacetimes can be decomposed into a Lorentzian product $\mathbb{R}\times\mathcal{M}$, where $\mathcal{M}$ is a Riemannian manifold describing the spatial structure of the spacetime and $\mathbb{R}$ is a real time dimension that characterizes causality  (Beem et al., 1996, Theorem 3.66 and 3.67). A point or event $(t,p)$ in the globally hyperbolic manifold is said to causally precede another event $(s,q)$ only if they are ordered $t\leq s$, where $t$ and $s$ represent time and $p$ and $q$ denote space coordinates, respectively. Unfortunately, simply using a single time dimension to define partial orders cannot encode most DAGs (see proof in Proposition 3), e.g. any directed tree with two or more leaves. Therefore, one must turn to richer causal structures (for instance with multiple time dimensions) to model greater directional complexity.
@@
The Neural Spacetime Model. We propose a trainable geometry with enough temporal and spatial dimensions to encode a broad class of weighted DAGs. In particular, the neural spacetime (NST) model utilizes a MultiLayer Perceptron (MLP) encoder that maps graph node features to an intermediate Euclidean latent space, which is then fragmented into space and time and processed in parallel by a neural metric network (inspired by the neural snowflake model from Borde & Kratsios (2023)) and a neural (partial) order. These embed nodes jointly as causally connected events in the continuous representation while respecting the one-hop distance (given by the DAG weights) and directionality of the original discrete graph structure, see Figure 1. The spatial component of our model, which considers $D$ spatial dimensions, leverages large-scale and asymptotic embedding approaches relevant in coarse and hyperbolic geometry (Nowak, 2005; Eskenazis et al., 2019), as well as the small-scale embedding-based approach of Borde & Kratsios (2023) inspired by fractal-type metric embeddings. These are often used to encode undirected graphs equipped with their (global) geodesic undirected distance (Bourgain, 1986; Matoušek, 1999; Gupta, 2000; Krauthgamer et al., 2004; Neiman, 2016; Elkin & Neiman, 2021; Andoni et al., 2018; Filtser, 2020; Kratsios et al., 2023a; Abraham et al., 2022), which implies that their local distance information can also be embedded (Abraham et al., 2007; Charikar et al., 2010). Note that the opposite is not true: optimizing for the local geometry does not guarantee a faithful representation of the global geometry (Ostrovska & Ostrovskii, 2019). On the other hand, the temporal component of our trainable geometry parameterizes an order structure $\lesssim^{\mathcal{T}}$ on multiple $T$ time dimensions. Thus, causality in our representation space simply means that a point $x\in\mathbb{R}^{D+T}$ precedes another $y\in\mathbb{R}^{D+T}$ in the sense that $x\lesssim^{\mathcal{T}}y$ (which only considers order in time $T$, see Section 3). When $T=1$, our geometry corresponds to using a Cauchy time function of the globally hyperbolic Lorentzian spacetimes to define a partial order (see Beem et al. (1996, Page 65) or Burtscher & Garcia-Heveling (2024)).
Contributions. Our contributions are as follows:
  • We propose a computationally tractable parametric construction of spacetimes (with multiple time dimensions) based on neural networks, ensuring that only valid geometries are learnable; hence, the term neural spacetimes (NSTs). Our approach decouples the representation into a product manifold, which models (quasi-)metrics and (partial) orders independently. It enables learning, via gradient descent, an isocausal embedding of DAGs (Section 3.1). We provide techniques to build, optimize, and stabilize artificial neural networks with fractalesque activation functions, upon which neural spacetimes are constructed (Section 3.3 and Appendix C).
  • Our main theoretical result, Theorem 1, provides a global embeddability guarantee showing that any (finite) poset can be embedded into a neural spacetime with a sufficient number of temporal and spatial dimensions, where the number of temporal dimensions equals the complexity (the so-called width of the poset) and the number of spatial dimensions depends on the desired embedding accuracy with respect to the shortest path distance on the underlying undirected graph. The embeddings are causal in time and asymptotically isometric in space. Furthermore, the neural spacetime model only requires $\mathcal{O}(k^{2})$ parameters to globally embed a weighted DAG with $k$ nodes. Secondly, Proposition 3 characterizes which posets can be embedded using at most two temporal dimensions. In particular, this shows that globally hyperbolic Lorentzian spacetimes are too rigid to capture the structure of many posets as they can only encode those with Hasse diagrams inducing a planar graph (see Example 2). Lastly, Theorem 2 shows that a broad class of posets can be embedded into a NST with at most two temporal dimensions. Here, the embeddings are causal in time and exactly isometric in space.
  • We experimentally validate our model on synthetic metric DAG embedding datasets, as well as real-world directed graphs such as the WebKB (Rozemberczki et al., 2021) and Dream5 (Marbach et al., 2012) datasets, which involve web hyperlink connections and gene expression networks, respectively. We compare neural spacetimes against other manifolds proposed in the spacetime representation learning literature and demonstrate the superior embedding capabilities of the proposed geometry.
In summary, in this work, we propose a parametric representation of spacetime geometries with embeddability guarantees for weighted DAGs, which can be practically implemented and optimized in a data-driven manner via gradient descent without requiring a priori a closed-form representation of the geometry.
@@ 2.1 Causal Structure
Directed Graphs. Consider a weighted directed graph, $G_{D}=(E_{D},V,W_{D})$, where $E_{D}$ represents a set of directed edges, $V$ a set of vertices, and $W_{D}$ the strength of the connections, respectively. We say that two vertices $u,v\in V$ are causally connected, denoted by $u\preccurlyeq v$, if there exists a sequence of directed edges, a path $\big{(}(\nu_{n},\nu_{n+1})\big{)}_{n=1}^{N-1}$ in $V$ with $\nu_{1}=u$ and $\nu_{N}=v$. Notice that causal connectivity is a transitive relation; that is, if $u\preccurlyeq v$ and $v\preccurlyeq w$, then $u\preccurlyeq w$, for each $u,v,w\in V$. The neighborhood $\mathcal{N}(u)$ of a node $u\in V$ is the set of nodes sharing an edge with $u$; i.e. $\mathcal{N}(u)\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\{v\in V:\,(u,v)\in E_{D}\mbox{ or }(v,u)\in E_{D}\}$. Every weighted directed graph $G_{D}=(E_{D},V,W_{D})$ is naturally a graph upon forgetting edge directions, meaning that it induces a weighted undirected graph $G=(E,V,W)$ where $E\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\{\{u,v\}\in V:\,(u,v)\in E_{D}\mbox{ or }(v,u)\in E_{D}\}$ and where $W$ is the symmetrization of $W_{D}$, $W(u,v)\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\max\{W_{D}(u,v),W_{D}(v,u)\}$.
@@
Posets. A broad class of logical relations can be encoded as directed graphs. We consider the class of partially ordered sets (posets): the pair or tuple $(V,\lesssim)$. In particular, a key characteristic of a poset is that for any two elements $u$ and $v$ in the set the following properties are satisfied: reflexivity, $\forall u\in V,u\preccurlyeq u$; antisymmetry, $u\preccurlyeq v\land v\preccurlyeq u\implies u=v$; and transitivity, $\forall u,v,w\in V,u\preccurlyeq v\land v\preccurlyeq w\implies u\preccurlyeq w$. Every poset can induce a directed graph structure on the vertex set $V$.
@@
The Hasse diagram of a poset $(V,\lesssim)$ is a directed graph $G_{D}=(E_{D},V)$ with $V\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}E_{D}$ and, for each $u,v\in V$, $(u,v)\in E_{D}$ iff 1) $x_{u}\leq x_{v}$, 2) $u\neq v$ and 3) there is no $w\in V\setminus\{u,v\}$ with $x_{u}\leq x_{w}\leq x_{v}$ (so $x_{u}\leq x_{v}$ is a “minimal relation”).
Our theoretical results in Section 3.2 concentrate on embedding guarantees for posets induced by DAGs, which are of interest in causal inference (Textor et al., 2016; Oates et al., 2016), causal optimal transport on DAGs (Eckstein & Cheridito, 2023), which is particularly important in sequential decisions making (Acciaio et al., 2020; Backhoff-Veraguas et al., 2020; Xu et al., 2020; Eckstein & Pammer, 2024; Kršek & Pammer, 2024; Gunasingam & Wong, 2024) and in robust finance (Bartl et al., 2021; Acciaio et al., 2024)), in applied category theory (Fong & Spivak, 2019) since any thin category is a poset (Chandler, 2019), and in biology applications such as gene regulatory networks (Marbach et al., 2012).

### v2

@@ Abstract
We propose a class of trainable deep learning-based geometries called Neural SpaceTimes (NSTs), which can universally represent nodes in weighted Directed Acyclic Graphs (DAGs) as events in a spacetime manifold. While most works in the literature focus on undirected graph representation learning or causality embedding separately, our differentiable geometry can encode both graph edge weights in its spatial dimensions and causality in the form of edge directionality in its temporal dimensions. We use a product manifold that combines a quasi-metric (for space) and a partial order (for time). NSTs are implemented as three neural networks trained in an end-to-end manner: an embedding network, which learns to optimize the location of nodes as events in the spacetime manifold, and two other networks that optimize the space and time geometries in parallel, which we call a neural (quasi-)metric and a neural partial order, respectively. The latter two networks leverage recent ideas at the intersection of fractal geometry and deep learning to shape the geometry of the representation space in a data-driven fashion, unlike other works in the literature that use fixed spacetime manifolds such as Minkowski space or De Sitter space to embed DAGs. Our main theoretical guarantee is a universal embedding theorem, showing that any $k$-point DAG can be embedded into an NST with $1+\mathcal{O}(\log(k))$ distortion while exactly preserving its causal structure. The total number of parameters defining the NST is sub-cubic in $k$ and linear in the width of the DAG. If the DAG has a planar Hasse diagram, this is improved to $\mathcal{O}(\log(k)+2)$ spatial and 2 temporal dimensions. We validate our framework computationally with synthetic weighted DAGs and real-world network embeddings; in both cases, the NSTs achieve lower embedding distortions than their counterparts using fixed spacetime geometries.
@@ 1 Introduction
footnotetext: Email addresses: chri6704@ox.ac.uk, kratsioa@mcmaster.ca, marcl@nvidia.com, xdong@robots.ox.ac.uk, michael.bronstein@cs.ox.ac.uk
@@
Graphs are a ubiquitous mathematical abstraction used in practically every branch of science from the analysis of social networks (Robins et al., 2007) to economic stability (Hurd et al., 2016) and genomics (Marbach et al., 2012). The breadth of social and physical phenomena encoded by large graphs has motivated the machine learning community to seek efficient representations (or embeddings) of graphs in continuous spaces. Since the structure of graphs typically has properties different from Euclidean geometry (e.g., trees exhibit exponential volume growth), a significant effort has been made to identify non-Euclidean geometries which can faithfully capture the structure of general graphs. Examples include hyperbolic embeddings of tree-like graphs (Ganea et al., 2018; Sonthalia & Gilbert, 2020; Shimizu et al., 2021; Kratsios et al., 2023a; b), spherical and toroidal embeddings of graphs with cycles (Schoenberg, 1942; Guella et al., 2016; Giovanni et al., 2022b), embeddings of graphs with several of these characteristics into product Riemannian geometries (Borde et al., 2023a; b; Giovanni et al., 2022a), combinations of constant curvature Riemannian manifolds (Lu et al., 2023) and manifolds with locally controllable Ricci curvature (Giovanni et al., 2022b).
@@
Figure 1: A Neural Spacetime (NST) is a learnable triplet $\mathcal{S}=(\mathcal{E},\mathcal{D},\mathcal{T})$, where $\mathcal{E}:\mathbb{R}^{N}\rightarrow\mathbb{R}^{D+T}$ is a (feature) encoder network, $\mathcal{D}:\mathbb{R}^{D+T}\times\mathbb{R}^{D+T}\to[0,\infty)$ is a learnable quasi-metric on $\mathbb{R}^{D}$ and $\mathcal{T}:\mathbb{R}^{D+T}\to\mathbb{R}^{T}$ is a learnable partial order on $\mathbb{R}^{T}$. Given an input Directed Acyclic Graph (DAG), $\mathcal{E}$ optimizes the location of the nodes $u,v,w$ as events in the spacetime manifold $\hat{u},\hat{v},\hat{w}$, while concurrently $\mathcal{D}$ and $\mathcal{T}$ learn the geometry of space and time themselves. The objective is to find a geometry that can faithfully represent, with minimal distortion, the metric geometry of the input DAG in space as well as its causal connectivity in time.
@@
The Directed Graph Embedding Problem. The description of many real-world systems requires directed graphs, for example, gene regulatory networks (Marbach et al., 2012), flow networks (Deleu et al., 2022), stochastic processes (Backhoff-Veraguas et al., 2020), or graph metanetworks (Lim et al., 2024) to name a few. Directed graphs (or digraphs) play a significant role in causal reasoning, where they are used to study the relationship between a cause and its effect in a complex system. These problems make use of representation spaces with causal structures, modeling the directional edges of a discrete graph as causal connections in a continuous representation space (Zheng et al., 2018). However, most approaches in causal representation learning primarily focus on capturing directional information while neglecting distance information. Recent works (Sim et al., 2021; Law & Lucas, 2023) have suggested that Lorentzian spacetimes from general relativity are suitable representation spaces for directed graphs, as these geometries possess an arrow of time that provides them with causal structure (see Appendix A.5) while also being capable of encoding distance.
Instead of considering a single arrow of time, in this paper we propose to turn to richer causal structures with multiple time dimensions, allowing us to model greater directional complexity.
Problem Formulation. Our goal is to learn a spacetime manifold representation where nodes in a DAG can be embedded as events. This should be done while capturing edge directionality in the form of causal structure (which includes modelling lack of connectivity as anti-chains), as well as the edge weights in the original discrete structure: these are represented as temporal ordering and spatial distance in the continuous embedding space.
@@
The Neural Spacetime Model. We propose a trainable geometry with enough temporal and spatial dimensions to encode a broad class of weighted DAGs. In particular, the Neural SpaceTime (NST) model utilizes a MultiLayer Perceptron (MLP) encoder that maps graph node features to an intermediate Euclidean latent space, which is then fragmented into space and time and processed in parallel by a neural metric network (inspired by the neural snowflake model from Borde & Kratsios (2023)) and a neural partial order. These embed nodes jointly as causally connected events in the continuous representation while respecting the one-hop distance (given by the DAG weights) and directionality of the original discrete graph structure, see Figure 1. The spatial component of our model, which considers $D$ spatial dimensions, leverages large-scale and asymptotic embedding approaches relevant in coarse and hyperbolic geometry (Nowak, 2005; Eskenazis et al., 2019), as well as the small-scale embedding-based approach of Borde & Kratsios (2023) inspired by fractal-type metric embeddings. These are often used to encode undirected graphs equipped with their (global) geodesic undirected distance (Bourgain, 1986; Matoušek, 1999; Gupta, 2000; Krauthgamer et al., 2004; Neiman, 2016; Elkin & Neiman, 2021; Andoni et al., 2018; Filtser, 2020; Kratsios et al., 2023a; Abraham et al., 2022), which implies that their local distance information can also be embedded (Abraham et al., 2007; Charikar et al., 2010). Note that the opposite is not true: optimizing for the local geometry does not guarantee a faithful representation of the global geometry (Ostrovska & Ostrovskii, 2019). On the other hand, the temporal component of our trainable geometry parameterizes an order structure $\lesssim^{\mathcal{T}}$ on multiple $T$ time dimensions. Thus, causality in our representation space simply means that a point $x\in\mathbb{R}^{D+T}$ precedes another $y\in\mathbb{R}^{D+T}$ in the sense that $x\lesssim^{\mathcal{T}}y$ (which only considers order in time $T$, see Section 3). When $T=1$, our geometry corresponds to using a Cauchy time function of the globally hyperbolic Lorentzian spacetimes to define a partial order (see Beem et al. (1996, Page 65) or Burtscher & Garcia-Heveling (2024)).
Our contributions are: (1) We propose a tractable neural network architecture for spacetimes (with multiple time dimensions), ensuring that only valid geometries are learnable; hence, the term neural spacetimes (NSTs). Our approach decouples the representation into a product manifold, which models (quasi-)metrics and (partial) orders independently. (2) Our main theoretical result, Theorem 1, provides a global embeddability guarantee showing that any (finite) poset can be embedded into a neural spacetime with a sufficient number of temporal and spatial dimensions. The embeddings are causal in time and asymptotically isometric in space. Furthermore, the neural spacetime model only requires $\mathcal{O}(k^{2})$ parameters to globally embed a weighted DAG with $k$ nodes. Theorem 2 shows that a broad class of posets can be embedded into a NST with at most two temporal dimensions. (3) We experimentally validate our model on synthetic metric DAG datasets, as well as real-world directed graphs that involve web hyperlink connections and gene expression networks, respectively.
@@ 2.1 Causal Structure
Directed Graphs. Consider a weighted directed graph, $G_{D}=(E_{D},V,W_{D})$, where $E_{D}$ represents a set of directed edges, $V$ a set of vertices, and $W_{D}$ the strength of the connections, respectively. We say that two vertices $u,v\in V$ are causally connected, denoted by $u\preccurlyeq v$, if there exists a sequence of directed edges, a path $\big{(}(\nu_{n},\nu_{n+1})\big{)}_{n=1}^{N-1}$ in $V$ with $\nu_{1}=u$ and $\nu_{N}=v$. Notice that causal connectivity is a transitive relation; that is, if $u\preccurlyeq v$ and $v\preccurlyeq w$, then $u\preccurlyeq w$, for each $u,v,w\in V$. The neighborhood $\mathcal{N}(u)$ of a node $u\in V$ is the set of nodes sharing an edge with $u$; i.e. $\mathcal{N}(u)\stackrel{{\scriptstyle\mbox{def.}}}{{=}}\{v\in V:\,(u,v)\in E_{D}\mbox{ or }(v,u)\in E_{D}\}$. Every weighted directed graph $G_{D}=(E_{D},V,W_{D})$ is naturally a graph upon forgetting edge directions, meaning that it induces a weighted undirected graph $G=(E,V,W)$ where $E\stackrel{{\scriptstyle\mbox{def.}}}{{=}}\{\{u,v\}\in V:\,(u,v)\in E_{D}\mbox{ or }(v,u)\in E_{D}\}$ and where $W$ is the symmetrization of $W_{D}$, $W(u,v)\stackrel{{\scriptstyle\mbox{def.}}}{{=}}\max\{W_{D}(u,v),W_{D}(v,u)\}$.
@@
Posets. A broad class of logical relations can be encoded as directed graphs. We consider the class of partially ordered sets (posets): the pair or tuple $(V,\lesssim)$. In particular, a key characteristic of a poset is that for any two elements $u$ and $v$ in the set the following properties are satisfied: reflexivity, $\forall u\in V,u\preccurlyeq u$; antisymmetry, $u\preccurlyeq v\land v\preccurlyeq u\implies u=v$; and transitivity, $\forall u,v,w\in V,u\preccurlyeq v\land v\preccurlyeq w\implies u\preccurlyeq w$. Every poset induces a DAG structure on the vertex set $V$.
@@
The Hasse diagram of a poset $(V,\lesssim)$ is a directed graph $G_{D}=(E_{D},V)$ with $V\stackrel{{\scriptstyle\mbox{def.}}}{{=}}E_{D}$ and $\forall u,v\in V$, $(u,v)\in E_{D}$ iff 1) $x_{u}\leq x_{v}$, 2) $u\neq v$ and 3) there is no $w\in V\setminus\{u,v\}$ with $x_{u}\leq x_{w}\leq x_{v}$ (so $x_{u}\leq x_{v}$ is a “minimal relation”).
Our theoretical results in Section 3.1 focus on embedding guarantees for posets induced by DAGs, which are of interest in causal inference (Textor et al., 2016; Oates et al., 2016), causal optimal transport on DAGs (Eckstein & Cheridito, 2023), which is particularly important in sequential decisions making (Acciaio et al., 2020; Backhoff-Veraguas et al., 2020; Xu et al., 2020; Eckstein & Pammer, 2024; Kršek & Pammer, 2024; Gunasingam & Wong, 2024) and in robust finance (Bartl et al., 2021; Acciaio et al., 2024)), in applied category theory (Fong & Spivak, 2019) since any thin category is a poset (Chandler, 2019), and in biology applications such as gene expressions (Marbach et al., 2012).